A bond has three years to​ maturity, a $ 2,000 face​ value, and a 6​% coupon rate with annual coupons. What is its yield to maturity if it is currently trading at $1,826​?
Answer #1

CY = C / P * 100, or CY = (B * CR / 100) / P

CY is the current yield,

C is the periodic coupon payment,

P is the price of a bond,

B is the par value or face value of a bond,

CR is the coupon rate.

P = C×(1 + r) -1 + C×(1 + r) -2 + . . . + C×(1 + r) -Y + B×(1 + r) -Y

The yearly coupon payment is $2000 × 6% = $120

1826=120*(1 + r) -1 + 120×(1 + r) -2  + 120×(1 + r) -3 + 2000×(1 + r) -3

Yield to Maturity: 9.47%